All production databases were patched successfully. If you experience any issues with Puget Sound systems, please contact TechHelp at (253) 879-8585.
The next production maintenance window is Sunday, November 29th, from 8:00 AM to noon. From 8:00 AM to noon, all database services may be intermittently or entirely unavailable, including: Cascade, FAMIS, Millennium, and Banner.
Updated 10/25/2009 @ noon — All production servers were patched successfully. If you experience any issues with Puget Sound systems, please contact TechHelp at (253) 879-8585.
The next production maintenance window is Sunday, October 25th, from 8:00 AM to noon.
From 8:00 AM to Noon, all services may be intermittently or entirely unavailable, including:
- Personal and departmental file shares (on Alexandria and Merlin2)
- Database Applications: Cascade, FAMIS, Millennium, and Banner
The following systems were patched today between 8am and noon:
- CBORDFSS (VENDOR DC)
- DM-1 to DM-8
- INTCHK (VM)
- KRONOS (VENDOR DC)
- LENEL (VENDOR Security)
- MICROS (VENDOR DC)
- ODYSSEY (VENDOR DC)
- VIAWARP (VM)
- VMMON (VM for now)
- CISCOWORKS (VM)
This month, on 10/15/2009 from 7am to 9pm we performed our monthly patch maintenance for PreProduction servers. We patched the following systems:
As of 9:00pm 10/13/09 all users have complete access to their email once again. Some messages into and out of these users accounts may be delayed as the system catches up.
Update as of 4:45PM 10/13/09
The outage of email for users EJ-LB continues as we run a database recovery process on the affected volume/users. The process could take up to 18 hours, completing by mid-morning Wednesday (10am).
This post will be updated as more information is available.
Exchange Email is unavailable for users with usernames starting with EJ-LB. Currently there is not a time estimate of when email for these users will be available again, but this post will be updated when the issue is corrected.
10/10/2009 The Aruba wirless controller was upgraded and additional new licensing applied starting at 4:30am. The controller was rebooted at 5:20am. The process is complete at 5:30am. Monitoring until 6:00am shows all access points are functioning properly.